# Self-Translate-Train: A Simple but Strong Baseline for Cross-lingual Transfer of Large Language Models

### v1

@@ Abstract
Cross-lingual transfer is a promising technique for utilizing data in a source language to improve performance in a target language. However, current techniques often require an external translation system or suffer from suboptimal performance due to over-reliance on cross-lingual generalization of multi-lingual pretrained language models. In this study, we propose a simple yet effective method called Self-Translate-Train. It leverages the translation capability of a large language model to generate synthetic training data in the target language and fine-tunes the model with its own generated data. We evaluate the proposed method on a wide range of tasks and show substantial performance gains across several non-English languages.
Self-Translate-Train: A Simple but Strong Baseline
for Cross-lingual Transfer of Large Language Models
@@ 1 Introduction
Cross-lingual transfer is a technique to solve tasks in a target language by leveraging training data from the other languages (Pikuliak et al., 2021). This has been increasingly feasible with the rise of multilingual pre-trained models, which are trained on multilingual corpora and capture commonalities across languages (Conneau et al., 2020; Xue et al., 2021; Scao et al., 2022). Capable multilingual models can perform tasks in a target language without being trained on task-specific data in that language, which is known as zero-shot cross-lingual transfer (Artetxe and Schwenk, 2019; Chen et al., 2021). This technique is expected to reduce the disparity between high-resource and low-resource languages.
Cross-lingual transfer is also exhibited in large language models (LLMs), which refers to auto-regressive language models with billion-scale parameters that are trained with a massive amount of text data (Brown et al., 2020; Touvron et al., 2023). A common approach for zero-shot cross-lingual transfer is fine-tuning the model with supervised training data available in a source languages, mostly English, and then applying the model to the target language (Chen et al., 2024; Shaham et al., 2024). However, we argue that this approach does not fully elicit the model’s cross-lingual capability as the model has no clue the input language at the test time. To achieve better cross-lingual performance, we let the model teach itself how to solve the task in the target language. In our proposed method, Self-Translate-Train, we produce the target language’s translation of the training data leveraging the strong capability of the LLM to generate text, and train the LLM with its own generated translation.
We evaluate Self-Translate-Train with several tasks including question answering, text-pair classification, and mathematical reasoning across multiple languages. Our experiments show that Self-Translate-Train consistently improves the performance of baselines given a multilingual capability of the LLMs. Our results indicate that we can achieve better cross-lingual performance by correctly elicit the model’s translation capability, which encourages further exploration of how to better utilize the model’s cross-lingual capability.
@@ 2.1 Cross-lingual Transfer Learning
There are two main approaches to transfer task knowledge across languages: data transfer and model transfer (Pikuliak et al., 2021).
Data transfer translates the source language data to the target language. In the Translate-test approach, models are trained on source language data and at inference time, the task inputs are translated into the source language (Conneau et al., 2018; Asai et al., 2018). Although the training stage is simple, it incurs additional translation costs at inference time. The Translate-train approach, on the other hand, translates the training data and the resulting model is used to predict the target language data directly (Conneau et al., 2018). Data transfer is quite effective in terms of performance (Hu et al., 2020), but one drawback is its requirement of additional translation systems.
Model transfer alleviates the need for translation systems by using multilingual pretrained models, which are trained on a large amount of data from multiple languages and capture the commonality between languages. These models can be fine-tuned on task-specific data in a single source language and generalize to solve the task in other languages (Pires et al., 2019; Mulcaire et al., 2019; Conneau et al., 2020), eliminating the need for translation systems.
Our approach, Self-Translate-Train, leverages LLMs’ translation and cross-lingual generalization capabilities. It combines the advantages of data transfer and model transfer by using explicit training signals in the target language while eliminating the need for external translation systems.
@@ 2.2 Self-Improvement of LLMs
LLMs have demonstrated remarkable text generation capabilities, which has been leveraged to generate training data for various purposes (Li et al., 2023b; Lee et al., 2024). The generated data can be used to further specialize the LLM itself for downstream applications, without requiring an extensive collection of additional data. This process can be viewed as a form of self-improvement (Bai et al., 2022; Huang et al., 2023; Sun et al., 2023; Li et al., 2023a).
Self-Translate-Train is also a self-improvement approach to specialize the LLM to a target language by translating the source language data to the target language.
@@ 5.1 Main Results
As the baseline, we fine-tune the LLM with the source language dataset . To ensure a fair comparison, we augment with the eight target language samples used for few-shot translation (Section 4.4). We then compare the baseline with the models fine-tuned on the data generated from Self-Translate-Train ( and ) in Table 3.
First, Self-Translate-Train is indeed an effective method; almost consistently outperforms the baseline . The only exception is Thai (th), where there is no significant improvement. This is likely due to the low translation quality of the model in Thai (Table 2).
The effectiveness of code-switching dataset is limited. When we add to , there is no significant improvement from adding only ( in Welch’s t-test). This indicates that the code-switching data does not provide additional information for the model to generalize in the task.
@@ 5.2 Does the model size matter?
The size of the language model can influence both its ability to generalize across languages and the quality of its translations, which in turn may impact the effectiveness of Self-Translate-Train. We compare the performance of Llama2 with different sizes, i.e., 7B, 13B, and 65B, on the math task in de, th, and zh (Figure 2).
The larger model size generally tends to perform better, and the improvement from Self-Translate-Train remains consistent across different model sizes. In Thai (th), we did not observe a significant improvement in the 7B model, but do in the larger models (13B and 65B), likely due to their better translation quality. The 7B model has a low Thai translation BLEU score of 1.9 (Table 2), while the 13B and 65B models have BLEU scores of 5.1 and 12.0, respectively.
The improvement in Thai (th) with the 70B model is the most significant (+19.8 average points). This implies that Self-Translate-Train is particularly effective when the model struggles with generalizing across the source and target languages but can still generate reasonable translations.
6 Conclusion
We introduced Self-Translate-Train, a method to improve cross-lingual transfer performance by generating synthetic training data in the target language. We validated its effectiveness on various tasks and languages, demonstrating substantial performance gains across several non-English languages. Self-Translate-Train is effective when the zero-shot cross-lingual transfer performance is suboptimal and the model can generate reasonable translations.
Self-Translate-Train neither requires external translation systems nor intensive additional data collection, making it a simple yet effective method for cross-lingual transfer. We encourage practitioners to try this approach as an improved baseline for cross-lingual transfer of LLMs.
Our research also shows that relying solely on the model’s generalization capability may be suboptimal, and there is a better way to elicit the cross-lingual capability of the model. We hope this work encourages further exploration of how to better utilize the model’s cross-lingual capability.
@@ 7 Limitations
Our method is based on the assumption that the model can generate reasonable translations in the target language. This may be challenging when the task inputs are long or complex. One solution is to split the input into smaller segments, as we have done with the SQuAD dataset (Section A.2).
Finally, when the task requires generating long and natural text, the quality of the generated translation matters more. If the translation quality is low, the model outputs may degrade due to translation errors or unnaturalness. The application of our method on more challenging tasks requires further investigation.
@@ A.2 Prompt Format for LLM Translation
The context passages in the SQuAD dataset are relatively long, and it is challenging for the LLM with a limited context window to fit the entire few-shot samples and the source text. To address this issue, we split the context into sentences using the spaCy library111https://spacy.io/ and translate them separately, i.e., the few-shot samples and the source text are sentences.
@@ A.4 Assessing the Translation Quality
To evaluate the translation quality, we use the BLEU score (Papineni et al., 2002) measured by the parallel data constructed from questions in the MGSM test set. The translation is performed in few-shot in-context learning with 8 translation samples constructed from the train set of the MGSM dataset. The BLEU score is calculated using the SacreBLEU library (Post, 2018)222https://github.com/mjpost/sacrebleu. As the tokenizer option, we use “13a” for de and ru, “flores101” for th, and “zh” for zh.
@@ A.5 Hyper-parameters for Fine-tuning
To increase the robustness of the results, we also conducted experiments with Qwen1.5-1.8B 333https://qwenlm.github.io/blog/qwen1.5/. While the model is mainly trained on Chinese and English data, it is also constructed with the multilingual use cases in mind.
| gsm8k | ||||||||||||
@@
Appendix C Frequently Asked Questions
In this section, we discuss questions that are outside the scope of the main topic of this paper but are somewhat relevant and may be of interest to readers.
@@ C.2 Does the synthetic data alone improve the performance in the target language?
C.3 Is tuning on generated by another model still effective?
Yes, if the other model can generate reasonable translations. Such approach can be seen as the Translate-train approach (Section 2.1) or sequence distillation from a teacher model (Kim and Rush, 2016).
As an upper-bound experiment using the math task, we fine-tune Llama2-7B on the synthetic data generated by gpt-3.5-turbo-0125 from the OpenAI API444https://openai.com/index/openai-api/, which produces high-quality translations across the languages explored in this paper (Table 6).
@@ C.3 Is tuning on generated by another model still effective?
Adding the synthetic data generated by gpt-3.5-turbo-0125 improves performance across languages.
However, the outputs from external models are often restricted in its usage555For example, Term of use of OpenAI API (January 31, 2024) restricts the usage of the outputs for training a model that competes with the API (https://openai.com/policies/terms-of-use/). Meta Llama 3 License (April 18, 2024) prohibits using the outputs to improve any other large language model (https://llama.meta.com/llama3/license/)., while the method explored in this paper can be used with the model at hand without other resources. Additionally, our interest in this paper is rather to explore the cross-lingual potential of the model itself and how to better utilize it.

### v2

# Self-Translate-Train: Enhancing Cross-Lingual Transfer of Large Language Models via Inherent Capability

@@ Abstract
Zero-shot cross-lingual transfer by fine-tuning multilingual pretrained models shows promise for low-resource languages, but often suffers from misalignment of internal representations between languages. We hypothesize that even when the model cannot generalize across languages effectively in fine-tuning, it still captures cross-lingual correspondence useful for cross-lingual transfer. We explore this hypothesis with Self-Translate-Train, a method that lets large language models (LLMs) to translate training data into the target language and fine-tunes the model on its own generated data. By demonstrating that Self-Translate-Train outperforms zero-shot transfer, we encourage further exploration of better methods to elicit cross-lingual capabilities of LLMs.
Self-Translate-Train: Enhancing Cross-Lingual Transfer
of Large Language Models via Inherent Capability
@@ 1 Introduction
Cross-lingual transfer is a technique to solve tasks in a target language by leveraging data from other languages (Pikuliak et al., 2021). A popular approach is to fine-tune a multilingual pre-trained model (Conneau et al., 2020; Xue et al., 2021; Scao et al., 2022) on a source language data (e.g., English) and then apply it to the target language. Capable models can solve tasks in the target language without seeing its data during fine-tuning, which is known as zero-shot cross-lingual transfer (Artetxe and Schwenk, 2019; Chen et al., 2021).
However, zero-shot cross-lingual transfer is often suboptimal, with a performance gap between source and target languages. One potential factor for this gap is representation misalignment between languages (Artetxe et al., 2023): the model does not capture the correspondence between languages effectively. We hypothesize that in some cases the model does capture the correspondence, but just cannot utilize it when fine-tuned only on the source language data; we need to elicit its cross-lingual capability in a different way.
To explore this hypothesis, we propose and investigate an approach called Self-Translate-Train, which leverages the powerful text generation capabilities of large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023). Instead of fine-tuning an LLM on source language data and hoping for generalization, Self-Translate-Train first let the model translate the training data into the target language, and fine-tunes the model on the data that the model itself generates. This process does not only rely on cross-lingual generalization that happens during fine-tuning, but also on the cross-lingual correspondence that the model captures and can be elicited through translation.
This approach is inspired by the Translate-Train (Conneau et al., 2018; Hu et al., 2020; Artetxe et al., 2023) approach, a method to translate training data to the target language, often with an external machine translation system. Notably, our findings reveal that the Translate-Train approach remains effective even without external resources.
We validate the effectiveness of Self-Translate-Train with three tasks (math, QA, NLI) across four target languages (de, ru, th, zh). The results show that better cross-lingual performance can be achieved by correctly eliciting the model’s translation capability, which encourages further exploration of a better method for cross-lingual transfer.
@@ 2.1 Cross-lingual Transfer Learning
There are two main approaches to transfer task knowledge across languages: model transfer and data transfer (Pikuliak et al., 2021).
Model transfer leverages multilingual pre-trained models. As the model is expected to capture the commonality between languages, when fine-tuned on source language data, it can generalize to solve the task in other languages (Pires et al., 2019; Mulcaire et al., 2019; Conneau et al., 2020).
Data transfer translates the source language data into the target language. In the Translate-Test approach, models are assumed to be specialized in the source language and the task inputs are translated into this language at test time (Conneau et al., 2018; Etxaniz et al., 2024). The Translate-Train approach translates the training data, and the trained model makes predictions in the target language (Conneau et al., 2018). Data transfer is often more effective than model transfer (Hu et al., 2020; Artetxe et al., 2023; Ebing and Glavaš, 2024).
Our approach, Self-Translate-Train, directly translates the training data as in data transfer, but also leverages the cross-lingual capabilities of the model itself as in model transfer.
@@ 2.2 Self-Improvement of LLMs
LLMs have remarkable text generation capabilities, which has been leveraged to generate training data for various purposes (Li et al., 2023b; Lee et al., 2024). The generated data can be used to further specialize the LLM itself for downstream applications, without requiring an extensive collection of additional data. This process can be viewed as a form of self-improvement (Bai et al., 2022; Huang et al., 2023; Sun et al., 2023; Li et al., 2023a).
Self-Translate-Train is also a self-improvement approach to specialize the LLM to a target language by translating the source language data.
@@ 5.1 Main Results
| FewShot |
As a baseline, we perform few-shot learning with the target language data with the pre-trained model (FewShot). We also fine-tune the LLM with the source language dataset . To ensure equal data resources, we augment with the eight target language samples used for few-shot translation (Section 4.4). We then compare these baselines with the models fine-tuned on the data generated from Self-Translate-Train ( and ) in Table 3111For experiments with other approaches, including translate-test and translation with external systems, see Appendix C..
Results show that fine-tuning on English () is more effective than few-shot learning in the target language (FewShot). This confirms that cross-lingual transfer from English is non-trivial and effective in this setting.
Notably, Self-Translate-Train is indeed an effective method; almost consistently outperforms the baseline . The only exception is Thai (th), where there is no significant improvement. This is likely due to the low translation quality of the model in Thai (Table 2).
However, the effectiveness of code-switching dataset is limited. When we add to , there is no significant improvement from adding only ( in Welch’s t-test). This indicates that the code-switching data does not provide additional information for the model to generalize in the task.
@@ 5.2 Does the model size matter?
The size of the LLM can influence both its ability to generalize across languages and the quality of its translations, which in turn may impact the effectiveness of Self-Translate-Train. We compare the performance of Llama2 with different sizes, i.e., 7B, 13B, and 65B, on the math task in de, th, and zh. For each model size, we generate with the corresponding model and then fine-tune itself.
Figure 2 shows that the larger model size generally tends to perform better, and the improvement from Self-Translate-Train remains consistent across different model sizes. In Thai (th), we did not observe a significant improvement in the 7B model, but do in the larger models (13B and 65B), likely due to their better translation quality. The 7B model has a low Thai translation BLEU score of 1.9 (Table 2), while the 13B and 65B models have BLEU scores of 5.1 and 12.0, respectively.
The improvement in Thai (th) with the 70B model is the most significant (+19.8 average points). This implies that Self-Translate-Train is particularly effective when the model struggles with generalizing across the source and target languages but can still generate their reasonable translations.
6 Discussion and Conclusion
We explored Self-Translate-Train, a method to improve cross-lingual transfer performance by generating synthetic training data in the target language. We validated its effectiveness on various tasks and languages, demonstrating substantial performance gains over baselines with the same data budget.
This result aligns with Etxaniz et al. (2024), who showed that the performance of few-shot learning can be improved by self-translating the task inputs. Our work demonstrates that self-translation by an LLM is also effective when applied to task-specific training data for fine-tuning.
These results suggest that we can improve the cross-lingual transfer performance of an LLM without any additional data, but by investing more computation to generate synthetic data. This finding is consistent with the recent trend of self-improvement methods for LLMs (e.g., Huang et al., 2023).
Future work could explore synthetic data generation for pre-training to improve cross-lingual capabilities of LLMs without intensive data collection. We hope that this work encourages further exploration of how to better elicit models’ cross-lingual capabilities.
@@ 7 Limitations
Our method is based on the assumption that the model can generate reasonable translations in the target language. However, for the low-resource target language or the datasets with out-of-domain text for the model, the translation quality can be not satisfactory for the Translate-Train approach (Ebrahimi and von der Wense, 2024).
Finally, when the task requires generating long and natural text, the quality of the generated translation matters more. If the translation quality is low, the model outputs may degrade due to translation errors or unnaturalness. The application of our method on more challenging tasks requires further technical considerations.
@@ A.2 Prompt Format for LLM Translation
The context passages in the SQuAD dataset are relatively long, and it is challenging for the LLM with a limited context window to fit the entire few-shot samples and the source text. To address this issue, we split the context into sentences using the spaCy library222https://spacy.io/ and translate them separately, i.e., the few-shot samples and the source text are sentences.
@@ A.4 Assessing the Translation Quality
To evaluate the translation quality, we use the BLEU score (Papineni et al., 2002) measured by the parallel data constructed from questions in the MGSM test set. The translation is performed in few-shot in-context learning with 8 translation samples constructed from the train set of the MGSM dataset. The BLEU score is calculated using the SacreBLEU library (Post, 2018)333https://github.com/mjpost/sacrebleu. As the tokenizer option, we use “13a” for de and ru, “flores101” for th, and “zh” for zh.
@@ A.5 Hyper-parameters for Fine-tuning
To increase the robustness of the results, we also conducted experiments with Qwen1.5-1.8B 444https://qwenlm.github.io/blog/qwen1.5/. While the model is mainly trained on Chinese and English data, it is also constructed with the multilingual use cases in mind.
| MGSM | ||||||||||||
@@
Appendix C Additional Exepriments
In this section, we discuss experiments that are outside the scope of the main topic of this paper but are somewhat relevant and may be of interest to readers.
@@ C.2 Does the synthetic data alone improve the performance in the target language?
C.3 What about the Translate-Test approach?
The translate-test approach, where non-English inputs are translated into English and fed into an English model (i.e., the model trained with ), can be an alternative to the trainslate-train approach. We perform experiments with the translate-test approach and provide additional context on the effectiveness of Self-Translate-Train. The tasks to use are the math and NLI tasks, whose outputs are language-independent 555We cannot simply apply the translate-test approach to QA, as the output language would be different from the original language. We could remedy this by further translating the output back to the original language (Asai et al., 2018), but this is out of the scope of our exploration here..
We first experiment with the Self-Translate-Test approach in few-shot learning settings (Etxaniz et al., 2024). This approach translates the test-time inputs to high-resource language (e.g., English) by few-shot translation and performs inference on the translated inputs.
We compare the Self-Translate-Test approach with a baseline without translation. The FewShot baseline uses 8-shot samples in the target language from the train set for inference on the test set. The Self-Translate-Test (few-shot) approach translates the test set to English and performs inference using English 8-shot samples. The results are shown in Table 11.
| MGSM | XNLI | |||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | |||||||||||||||||
| FewShot |
| Self-Translate-Test (few-shot) |
We can see that the Self-Translate-Test approach is effective both in the math and NLI tasks, which conforms to the results in Etxaniz et al. (2024).
Next, we validate the effectiveness in fine-tuning settings. Specifically, we feed the model fine-tuned with English task-specific data with test inputs translated into English by few-shot translation of the pre-trained model. The results are shown in Table 12.
| MGSM | XNLI | |||||||||||||||||||||||
| de | ru | th | zh | de | ru | th | zh | |||||||||||||||||
| + Self-Translate-Test |
We observe that the Self-Translate-Test approach remains effective when combined with a fine-tuned model. This strengthen the argument that the self-translate approach is an effective way to improve the cross-lingual performance of the model under the same data budget. However, this approach is not always practical, as it requires both the pre-trained model for few-shot translation and the fine-tuned model to be available during deployment.
C.4 How effective is generated by an external system?
When the synthetic training data is generated by an external system, such approach can be seen as the Translate-Train approach (Section 2.1) or sequence distillation from a teacher model (Kim and Rush, 2016). As an upper-bound experiment using the math task, we fine-tune Llama2-7B on the synthetic data generated by gpt-3.5-turbo-0125 from the OpenAI API666https://openai.com/index/openai-api/, which produces high-quality translations across the languages explored in this paper (Table 6).
@@ C.4 How effective is generated by an external system?
| from Llama2-7B |
| from Llama2-7B |
| from gpt-3.5-turbo-0125 |
| from gpt-3.5-turbo-0125 |
The synthetic data generated by gpt-3.5-turbo-0125 improves performance across languages over the baseline and the synthetic data generated by Llama2-7B. This confirms that the quality of the synthetic data is crucial for the performance improvement.
Note that the outputs from external models are often restricted in its usage777For example, Term of use of OpenAI API (January 31, 2024) restricts the usage of the outputs for training a model that competes with the API (https://openai.com/policies/terms-of-use/). Meta Llama 3 License (April 18, 2024) prohibits using the outputs to improve any other large language model (https://llama.meta.com/llama3/license/)., while the method explored in this paper can be used with the model at hand without other resources. Additionally, our interest in this paper is rather to explore the cross-lingual potential of the model itself and how to better utilize it.